# CangLing-KnowFlow: A Unified Knowledge-and-Flow-fused Agent for Comprehensive Remote Sensing Applications

### v1

@@ 4.3 Evaluation Metrics
To comprehensively assess agent performance, we designed a multi-faceted evaluation protocol that examines both the final outcome of a task and the procedural fidelity, logical coherence, and step-wise accuracy of the execution trace. Our protocol is divided into two primary categories: end-to-end evaluation and step-wise execution evaluation.
@@ 4.3.1 End-to-End Evaluation
This category assesses the overall effectiveness and efficiency of an agent in accomplishing a given RS task from start to finish. It is used to compare the performance of different agentic frameworks (e.g., ReAct, Reflexion, CangLing-KnowFlow) across our benchmarks.
@@
Planning Efficiency: This set of metrics quantifies the cost and autonomy of the agent’s planning process.
Number of Tool Calls (NTC): The total count of tool invocations required to complete the task. A smaller number generally indicates a more efficient plan.
Number of Interactions (NI): The frequency of required user interventions or confirmations. A lower count signifies greater autonomy.
@@ 4.3.2 Step-wise Execution Evaluation
This category provides a fine-grained analysis of the procedural correctness of our CangLing-KnowFlow agent’s execution trace. To achieve this, we adopt the fine-grained, step-wise evaluation metrics proposed in the GTA benchmark [wang2024gta]. This methodology is designed to assess the procedural fidelity of an agent’s reasoning and action trace at each step of the workflow, offering deep insights into its operational capabilities.
Instruction Following Accuracy (InstAcc): This metric evaluates whether the agent’s output at each step adheres strictly to the predefined format (e.g., Thought: … Action: …). It serves as a fundamental check of the model’s ability to generate structured output.
Tool Selection Accuracy (ToolAcc): This measures whether the tool selected by the agent at a given step matches the one in the ground-truth workflow. It is a core metric for evaluating the agent’s task decomposition and planning abilities.
@@ 5.1 Experimental Setup
To rigorously evaluate the performance and efficacy of our proposed CangLing-KnowFlow framework, we designed a comprehensive experimental protocol. This section details the evaluation benchmarks, the baseline methods against which our framework is compared, the suite of LLMs used as backbones, and the specific metrics employed for quantitative assessment.
Benchmarks. The evaluation was conducted on two complementary benchmarks: the newly developed KnowFlow-Bench and the public ThinkGeo benchmark. The KnowFlow-Bench is engineered specifically to assess an agent’s end-to-end capabilities in orchestrating complex, multi-step RS workflows, from high-level task planning and tool invocation to dynamic, adjustments in response to execution failures. To further test the generalization capabilities and adaptability of our framework, we conducted evaluations on ThinkGeo. This allowed us to assess CangLing-KnowFlow’s performance in a novel environment with a distinct set of predefined tools and task structures.
Baseline Methods. The performance of CangLing-KnowFlow was systematically benchmarked against three representative agent paradigms, establishing a clear spectrum of capability:
@@
    ReAct [yao2023react]: A foundational paradigm that synergizes reasoning and acting by generating interleaved “Thought” and “Action” traces. It serves as the standard baseline for structured tool use. In our experiments, the ReAct agent operates using standard prompting strategies while accessing the identical tool library, providing a benchmark for performance without external procedural knowledge augmentation.
@@
    Reflexion [shinn2023reflexion]: An advanced evolution of ReAct that incorporates a verbal reinforcement learning mechanism. Upon detecting a failure, the agent analyzes its execution trace to generate textual self-reflections, which act as heuristic feedback for subsequent trials. While equipped with the same tools as the other agents, this baseline enables us to assess the efficacy of purely linguistic self-correction compared to our structured, graph-based evolutionary memory.
@@
Backbone LLMs. To ensure that our findings are robust and not contingent on a single model’s idiosyncrasies, all experiments were replicated across a diverse and powerful set of state-of-the-art LLMs. This selection, detailed in Table 1, includes leading open-source models (e.g., Llama3-70B [dubey2024llama], Claude 3.7 Sonnet) and top-tier commercial models (e.g., GPT-4 [achiam2023gpt]), spanning a wide range of architectures and capabilities.
Evaluation Metrics. We employed the comprehensive evaluation protocol defined in Section 4.3 to provide a holistic assessment of agent performance.
@@ 5.2 Overall Performance Comparison
To validate the effectivity of our framework empirically, we conducted a comprehensive comparison against established baseline methods on our proposed KnowFlow-Bench. The results, presented systematically in Table 1, offer a clear and consistent tendency– the CangLing-KnowFlow framework demonstrates a significant improvement across all evaluated metrics and LLM backbones.
@@
Mistral-7B [jiang2023diego] 2023 Open TSR (%) ↑ 54.6 (38.9) 58.9 (43.6) 60.2 (45.1) 68.4 (54.2)
@@
Llama3-8B [dubey2024llama] 2024 Open TSR (%) ↑ 58.3 (42.1) 62.1 (46.8) 63.7 (48.2) 71.5 (57.9)
@@
Llama3-70B [dubey2024llama] 2024 Open TSR (%) ↑ 73.4 (60.4) 76.2 (64.2) 77.1 (65.1) 83.6 (74.6)
@@
DeepSeek-R1 [guo2025deepseek, guo2025deepseek_arxiv] 2025 Open TSR (%) ↑ 83.4 (77.4) 86.9 (82.9) 87.8 (84.8) 93.1 (89.1)
@@
The most critical metric, Task Success Rate, reveals the profound impact of our architecture. For both simple and complex tasks, CangLing-KnowFlow consistently achieves the highest success rates, often by a significant margin. For example, when powered by the Claude 4 Sonnet model, our framework boosts the TSR on complex tasks to 95.1%, a notable improvement over the 90.6% achieved by the next-best method, Reflexion. This underscores the efficacy of integrating an expert-informed PKB, which provides a scientifically valid basis for planning and fundamentally mitigates the planning fallacies endemic to less-constrained agent architectures.
Furthermore, the results highlight the great potential for improving planning efficiency and autonomy. The FPA of CangLing-KnowFlow is consistently higher than all baselines, indicating that the initial workflows derived from our knowledge base are of superior quality and less prone to immediate failure. Consequently, this leads to a substantial reduction in both the NTC and the NI. By providing a more direct and validated path to the solution, our framework minimizes the inefficient trial-and-error cycles characteristic of ReAct-based agents, thereby enhancing automation and reducing computational overhead.
@@
To assess the generalizability and robustness of our framework beyond its native environment, we also evaluated its performance on the public ThinkGeo benchmark [shabbir2025thinkgeo]. We consider this to be the most critical validation of our framework. It serves as the definitive test of whether the agent’s capabilities are intrinsic to its architecture or merely artifacts of overfitting to the specific tools and tasks of its native environment. By confronting the agent with ThinkGeo’s distinct set of 14 tools and unique task distribution, we rigorously evaluated its ability to adapt its structured reasoning and dynamic adjustment mechanisms to a completely novel and uncertain problem space.
We selected several representative LLM backbones and compared our framework’s performance against the baselines. The results of this cross-benchmark evaluation are presented in Table 2. The data demonstrate that even when operating within an unfamiliar toolset, the core principles of CangLing-KnowFlow—leveraging a structured workflow guided by knowledge and dynamic adjustment—continue to yield a significant performance uplift, confirming the architectural soundness and transferability of our approach.
@@ 5.3 Ablation Study
To deconstruct the architectural contributions of our framework and ascertain the individual importance of its core components, we conducted a comprehensive ablation study. Using the high-performing Claude 4 Sonnet as the backbone LLM, we systematically disabled each of the three foundational mechanisms of CangLing-KnowFlow: the expert-informed Workflow Library (WL), the Dynamic Adjustment (DA) module, and the experience-driven Learning Capability (LC). The quantitative results, detailed in Table 3, unequivocally demonstrate that each component is integral to the framework’s synergistic performance.
@@
  • 1.
    The Indispensable Role of the Workflow Library (w/o WL): The most substantial performance degradation was observed upon the removal of the Workflow Library. In this configuration, the agent, stripped of its expert-guided procedural knowledge, reverted to a purely LLM-driven, unconstrained planning mode. This resulted in a dramatic decline in the TSR to 88.7% and a fall in FPA to 67.1%. Notably, planning efficiency deteriorated significantly, as evidenced by a marked increase in the NTC to 13.2 and NI to 4.3. This outcome robustly validates our central hypothesis: an expert-informed knowledge base is the foundational basis that prevents planning hallucination and constrains the agent to methodologically sound and efficient operational pathways.
  • 2.
    The Critical Function of Dynamic Adjustment (w/o DA): Disabling the Dynamic Adjustment module, which handles runtime failures, provided critical insights into the framework’s robustness. While the FPA remained high at 75.9%, reflecting the quality of the initial plans derived from the WL, the final TSR decreased to 91.5%. This gap between initial plan quality and final success underscores a crucial reality of real-world applications: even well-structured plans are vulnerable to failures stemming from data variability or environmental stochasticity. The DA module serves as the essential safety net, endowing the agent with the resilience needed to navigate these unforeseen challenges and ensure task completion. Its absence renders the agent brittle, unable to recover from otherwise manageable errors.
  • 3.
    The Optimizing Effect of Learning Capability (w/o LC): The removal of the Learning Capability module, which allows the agent to evolve from experience, resulted in a discernible, albeit less dramatic, decline in accuracy (TSR of 92.8%). The LC functions as a long-term optimization engine. By solidifying validated ad-hoc repairs into new workflow templates and attributing failures to heuristic rules, it continuously refines the agent’s knowledge base. Its absence does not severely impair the agent’s immediate problem-solving ability, but rather curtails its potential for growth, preventing it from becoming more efficient and autonomous over time. This suggests that the utility of the LC module is only partially evaluated in this study; we anticipate that its value would be significantly amplified in longer-term implementations where the benefits of cumulative learning can accrue. This component is key to transitioning the agent from a static tool to an evolving, intelligent system.
In summary, the ablation study confirms that the superior performance of CangLing-KnowFlow is not due to any single feature but to the tightly integrated synergy of its components. The Workflow Library provides the essential strategic guidance, the Dynamic Adjustment module ensures tactical robustness, and the Learning Capability enables long-term evolution.
@@ 5.4 LLM’s Intrinsic Capability
As evidenced in Table 4, there is a large positive correlation between an LLM’s inherent abilities–such as instruction following (Inst Acc.), tool selection (Tool Sel. F1), and parameter accuracy (Param. Acc.)–and its potential as an agent backbone. Top-tier models like Claude 4 Sonnet and GPT-5 demonstrate superior foundational skills. However, a myopic focus on the LLM alone would be misleading. The more profound insight emerges from cross-referencing these intrinsic capabilities with the end-to-end task performance in Table 1. For example, Llama3-70B [dubey2024llama], a model with respectable but not top-tier intrinsic scores, achieves a TSR of 83.6% (simple) and 74.6% (complex) when empowered by our CangLing-KnowFlow framework. It is notably better to that of a far more powerful model, GPT-4 [achiam2023gpt], when constrained to the ReAct framework (TSR of 82.9% / 74.9%). This demonstrates that the CangLing-KnowFlow architecture does not merely facilitate the LLM’s existing abilities; it fundamentally enhances and guides them, enabling a moderately capable model to outperform a superior one that lacks proper guidance.
@@ 6 Conclusion
Empirical validations across diverse benchmarks confirm the efficacy and robustness of our approach. On the domain-specific KnowFlow-Bench, the framework (powered by Claude 4 Sonnet) achieved peak Task Success Rates of 96.1% and 95.1% on simple and complex tasks respectively, significantly outperforming ReAct baselines while reducing tool calls by approximately 38%. Crucially, this reasoning capability was found to be transferable: evaluations on the generalization ThinkGeo benchmark demonstrated a performance doubling (20.3% vs. 9.0%) against baselines in unseen environments. Furthermore, ablation studies quantitatively identified the expert-informed Workflow Library as the foundational element for preventing planning hallucinations, with its removal causing the most significant degradation in both accuracy and efficiency.
By endowing agents with the cognitive capacities to reason, adapt, and learn within complex, dynamic environments, rather than merely executing static procedures, we establish a new paradigm for efficient, reliable, and scalable discovery. Looking ahead, our research will advance this paradigm along several key frontiers. We will focus on expanding the PKB through semi-automated knowledge acquisition and engineering multi-agent collaborative workflows for highly complex scientific campaigns. More profoundly, we will direct CangLing-KnowFlow’s core principles toward accelerating mechanism-driven discovery in new domains where procedural integrity is paramount. A primary objective is to harness this framework to unravel the complex, interdisciplinary processes central to fields such as global sustainability and ecology, where RS science is poised to unlock fundamental new insights [ghamisi2025geospatial]. This trajectory positions CangLing-KnowFlow not merely as an observation tool, but as a proactive partner in scientific inquiry.

### v2

@@ 4.3 Evaluation Metrics
To ensure a holistic assessment of agent performance, we designed a comprehensive evaluation protocol that rigorously evaluates both the final task outcome and the procedural fidelity, logical coherence, and step-wise accuracy of the execution trace. Accordingly, our protocol is organized into two primary categories, end-to-end outcome evaluation and step-wise process evaluation.
@@ 4.3.1 End-to-End Evaluation
The first category assesses the overall effectiveness and efficiency of an agent in accomplishing a given RS task from start to finish. It is used to compare the performance of different agentic frameworks (e.g., ReAct, Reflexion, CangLing-KnowFlow) across our benchmarks.
@@
Planning Efficiency: This set of metrics quantifies the cost and autonomy of the agent’s planning process, primarily consists of (1) Number of Tool Calls (NTC): The total count of tool invocations required to complete the task. A smaller number generally indicates a more efficient plan, and (2) Number of Interactions (NI): The frequency of required user interventions or confirmations. A lower count signifies greater autonomy.
@@ 4.3.2 Step-wise Execution Evaluation
The second category provides a fine-grained analysis of the procedural correctness of our CangLing-KnowFlow agent’s execution trace. To achieve this, we adopt the fine-grained, step-wise evaluation metrics proposed in the GTA benchmark [wang2024gta]. This manner is designed to assess the procedural fidelity of an agent’s reasoning and action trace at each step of the workflow, offering deep insights into its operational capabilities.
Instruction Following Accuracy (InstAcc): This metric evaluates whether the agent’s output at each step follows strictly to the predefined format (e.g., Thought: … Action: …). It serves as a fundamental check of the model’s ability to generate structured output.
Tool Selection Accuracy (ToolAcc): It measures whether the tool selected by the agent at a given step matches the one in the ground-truth workflow, which is a core metric for evaluating the agent’s task decomposition and planning abilities.
@@ 5.1 Experimental Setup
To rigorously evaluate the performance and efficacy of our proposed CangLing-KnowFlow framework, we designed a comprehensive experimental protocol. This section details the evaluation benchmarks, the selected baseline methods for comparison, the suite of LLMs used as backbones, and the specific metrics employed for quantitative assessment.
As discussed before, the evaluation was conducted on two complementary benchmarks: our developed KnowFlow-Bench and the public ThinkGeo benchmark [shabbir2025thinkgeo]. The KnowFlow-Bench is tailored to assess an agent’s end-to-end capabilities in orchestrating complex, multi-step RS workflows, from high-level task planning and tool invocation to dynamic adjustments in response to execution failures. The ThinkGeo is used to further test the generalization capabilities and adaptability of our framework, allowing us to assess CangLing-KnowFlow’s performance in a novel environment with a distinct set of predefined tools and task structures.
Given the lack of competitors in this emerging field, we established the following settings as baselines for comparison:
@@
    ReAct [yao2023react]: A foundational paradigm that synergizes reasoning and acting by generating interleaved “Thought” and “Action” traces. It serves as the standard baseline, employing standard prompting strategies and the shared toolset. This setup establishes performance metrics in the absence of external procedural knowledge augmentation.
@@
    Reflexion [shinn2023reflexion]: An advanced evolution of ReAct that incorporates a verbal reinforcement learning mechanism. When a failure is detected, the agent can analyze its execution trace to generate textual self-reflections, which act as heuristic feedback for subsequent trials. While equipped with the same tools as the other agents, this baseline enables us to assess the efficacy of purely linguistic self-correction compared to our structured, graph-based evolutionary memory.
@@
To ensure the robustness of our analysis and mitigate potential biases inherent to specific architectures, we evaluated our framework across a diverse family of state-of-the-art LLMs. As detailed in Table 1, our selection encompasses both leading proprietary models (e.g., GPT-4 [achiam2023gpt], Claude 3.7 Sonnet) and high-performance open-weight models (e.g., Llama3-70B [grattafiori2024llama]). All evaluations strictly follows the comprehensive protocol defined in Section 4.3 to guarantee a holistic assessment.
@@ 5.2 Overall Performance Comparison
To validate the effectivity of our framework empirically, we conducted a comprehensive comparison against established baseline methods on our proposed KnowFlow-Bench. The detailed results summarized in Table 1 offer a clear and consistent tendency that the CangLing-KnowFlow framework wins a significant improvement across all evaluated metrics and LLM backbones.
@@
Mistral-7B [jiang2023mistral] 2023 Open TSR (%) ↑ 54.6 (38.9) 58.9 (43.6) 60.2 (45.1) 68.4 (54.2)
@@
Llama3-8B [grattafiori2024llama] 2024 Open TSR (%) ↑ 58.3 (42.1) 62.1 (46.8) 63.7 (48.2) 71.5 (57.9)
@@
Llama3-70B [grattafiori2024llama] 2024 Open TSR (%) ↑ 73.4 (60.4) 76.2 (64.2) 77.1 (65.1) 83.6 (74.6)
@@
DeepSeek-R1 [guo2025deepseek] 2025 Open TSR (%) ↑ 83.4 (77.4) 86.9 (82.9) 87.8 (84.8) 93.1 (89.1)
@@
Most notably, the TSR metric highlights the distinct advantage of our proposed approach. Across both simple and complex tasks, CangLing-KnowFlow consistently achieves the highest success rates, outperforming baselines by significant margins. For example, when powered by the Claude 4 Sonnet model, our framework boosts the TSR on complex tasks to 95.1%, a substantial improvement over the 90.6% achieved by the nearest competitor, Reflexion. This result validates the efficacy of integrating an expert-informed PKB, which guides the agent with scientifically valid prior knowledge, thereby effectively mitigating the logical hallucinations and inefficient planning often observed in unconstrained agent architectures.
Furthermore, the results highlight the framework’s capability to significantly enhance planning efficiency and autonomy. CangLing-KnowFlow consistently yields higher FPA scores than all baselines, indicating that initial workflows derived from our knowledge base are inherently robust and less susceptible to early-stage failure. Consequently, this leads to a substantial reduction in both the NTC and the NI. By offering a validated, direct path to the solution, our approach curtails the redundant iterative loops that are typical in ReAct-based agents, thereby lowering computational costs and enhancing the degree of automation.
@@
To verify the cross-domain generalizability of our framework, we extended our evaluation to the public ThinkGeo benchmark [shabbir2025thinkgeo]. We consider this experiment a pivotal validation step. It serves as a crucial test to determine whether the agent’s capabilities are intrinsic to its architecture rather than merely artifacts of overfitting to specific internal tools and tasks. By subjecting the agent to ThinkGeo’s distinct inventory of 14 tools and diverse tasks, we strictly evaluated its capacity to adopt its structured reasoning and dynamic adjustment mechanisms to an unseen case.
We benchmarked our framework against baselines using a selection of representative LLM backbones. The results, detailed in Table 2, reveal a compelling trend: even when operating within an unfamiliar toolset, CangLing-KnowFlow maintains a significant performance advantage. This demonstrates that our core, i.e., leveraging knowledge-guided structured workflows and dynamic adjustments, remains effective regardless of the specific tool environment, thereby confirming the architectural soundness and high transferability of our approach.
@@ 5.3 Ablation Study
To quantify the individual contributions of our framework’s core components, we conducted an ablation study using the Claude 4 Sonnet backbone. We systematically disabled the Workflow Library (WL), Dynamic Adjustment (DA), and Learning Capability (LC). As shown in Table 3, the results confirm that each module is indispensable to the system’s overall performance.
@@
Specifically, the most substantial performance degradation occured upon the removal of the Workflow Library. In this configuration, the agent, stripped of its expert-guided procedural knowledge, reverted to a purely LLM-driven, unconstrained planning mode. This led to a dramatic decline in the TSR to 88.7% and a sharp drop in FPA to 67.1%. Notably, planning efficiency deteriorated significantly, as evidenced by by the surge in NTC to 13.2 and NI to 4.3. This outcome validates our central hypothesis, an expert-informed knowledge base is the foundational basis that prevents planning hallucination, constraining the agent to operational pathways that are both methodologically sound and computationally efficient.
Conversely, disabling the Dynamic Adjustment module provided critical insights into the framework’s operational robustness. While the FPA remained relatively high at 75.9%, reflecting the quality of the initial plans derived from the WL, the final TSR decreased to 91.5%. This divergence between initial plan quality and final success highlights a crucial reality of real-world applications: even well-structured plans are vulnerable to failures stemming from data variability or environmental stochasticity. The DA module acts as the essential safety net, endowing the agent with the resilience required to navigate these unforeseen challenges and ensure task completion. Its absence renders the agent brittle, leaving it unable to recover from otherwise manageable execution errors.
Finally, the removal of the Learning Capability module resulted in a moderate yet discernible decline in accuracy (TSR of 92.8%). The LC functions as a long-term optimization engine. By codifying validated ad-hoc repairs into new workflow templates and attributing failures to heuristic rules, it continuously refines the agent’s knowledge base. While its absence does not severely impair immediate problem-solving abilities, it restricts the evolutionary potential of the agent, preventing gains in efficiency and autonomy over time. This suggests that the utility of the LC module is underestimated in this short-term study; we anticipate its value would be significantly amplified in long-term implementations where the benefits of cumulative learning can fully accrue. This component is key to transitioning the agent from a static tool to an evolving, intelligent system.
In summary, the ablation study confirms that the superior performance of our CangLing-KnowFlow stems not from any single feature, but from the orchestrated synergy of its components. The Workflow Library provides essential strategic guidance, the Dynamic Adjustment module ensures tactical robustness, and the Learning Capability enables long-term evolution.
@@ 5.4 LLM’s Intrinsic Capability
As evidenced in Table 4, there is a large positive correlation between an LLM’s inherent abilities–such as instruction following (Inst Acc.), tool selection (Tool Sel. F1), and parameter accuracy (Param. Acc.)–and its potential as an agent backbone. Top-tier models like Claude 4 Sonnet and GPT-5 demonstrate superior foundational skills. However, a myopic focus on the LLM alone would be misleading. The more profound insight emerges from cross-referencing these intrinsic capabilities with the end-to-end task performance in Table 1. For example, Llama3-70B [grattafiori2024llama], a model with respectable but not top-tier intrinsic scores, achieves a TSR of 83.6% (simple) and 74.6% (complex) when empowered by our CangLing-KnowFlow framework. It is notably better to that of a far more powerful model, GPT-4 [achiam2023gpt], when constrained to the ReAct framework (TSR of 82.9% / 74.9%). This demonstrates that the CangLing-KnowFlow architecture does not merely facilitate the LLM’s existing abilities; it fundamentally enhances and guides them, enabling a moderately capable model to outperform a superior one that lacks proper guidance.
@@ 6 Conclusion
Empirical validations across diverse benchmarks confirm the efficacy and robustness of our approach. On the domain-specific KnowFlow-Bench, the framework (powered by Claude 4 Sonnet) achieved peak Task Success Rates of 96.1% and 95.1% on simple and complex tasks, respectively, significantly outperforming ReAct baselines while reducing tool calls by approximately 38%. We observed a similar trend on the ThinkGeo benchmark, where our method achieved more than double the baseline performance, increasing from 9.0% to 20.3% in unseen environments. Furthermore, ablation studies quantitatively identified the expert-informed Workflow Library as the foundational element for preventing planning hallucinations, with its removal causing the most significant degradation in both accuracy and efficiency.
By endowing agents with the cognitive capacities to reason, adapt, and learn within complex, dynamic environments, rather than merely executing static procedures, we establish a new paradigm for efficient, reliable, and scalable discovery. Looking ahead, our research will advance this paradigm along several key frontiers. We will focus on expanding the PKB through semi-automated knowledge acquisition and engineering multi-agent collaborative workflows for highly complex scientific campaigns. More profoundly, we will direct CangLing-KnowFlow’s core principles toward accelerating mechanism-driven scientific discovery in new domains where procedural integrity is paramount. For example, a primary objective can be harnessing this framework to unravel the complex, interdisciplinary processes central to fields such as global sustainability and ecology, where RS science is poised to unlock fundamental new insights. This will position CangLing-KnowFlow not merely as an observation tool, but as a proactive partner in scientific inquiry.